work out the value of the series
1+2+4+5+7+8+10+11+......+50

Answer by richard1234(7193)

1+2+3+...+50 = 50*51/2 = 1275

We are subtracting off the multiples of 3 from 3 to 48, inclusive, in which 3+6+9+...+48 = 3(1+2+...+16) = 3(16*17/2) = 408. The sum is equal to 1275 - 408 = 867.